# LlaSMol: Advancing Large Language Models for Chemistry with a Large-Scale, Comprehensive, High-Quality Instruction Tuning Dataset

### v1

@@ 3 Preliminaries
Molecules can be represented in multiple ways, such as SMILES strings, IUPAC names, and molecular formulas. SMILES strings use a sequence of symbols to encode the 2D structures of molecules (Weininger, 1988). A molecule can have multiple SMILES strings; a canonical SMILES for the molecule is unique and deterministic. For example, the canonical SMILES representation of glucose is C(C1C(C(C(C(O1)O)O)O)O)O. Molecular formulas represent a molecule by enumerating the type and number of atoms in the molecule (Solomons et al., 2022). For example, the molecular formula for glucose is . IUPAC names are formal names based on natural language elements, which follow the systematic rules set by the International Union of Preferred and Applied Chemistry (IUPAC) (Favre & Powell, 2014). These names are derived from the structures and functional groups of molecules, and are intended to be human-readable. For example, the IUPAC name for glucose is (3R,4S,5S,6R)-6-(hydroxymethyl)oxane-2,3,4,5-tetrol.
@@
4 SMolInstruct
4.1 Overview of SMolInstruct
@@
SMolInstruct is a large-scale instruction tuning dataset that centers around small molecules. It contains a total of 14 chemistry tasks, as illustrated in Figure 1.
(1) We include four name conversion tasks, namely converting IUPAC to molecular formula (NC-I2F), converting IUPAC to SMILES (NC-I2S), converting SMILES to molecular formula (NC-S2F), and converting SMILES to IUPAC (NC-S2I). They are designed to enable deep understanding of molecular structures and representations, which should serve as the fundamental knowledge for chemistry LLMs.
(2) Additionally, six property prediction tasks are integrated, including PP-ESOL for water solubility (Mobley & Guthrie, 2014), PP-Lipo for octanol/water distribution coefficient (Poole & Poole, 2003), PP-BBBP for blood-brain barrier penetration (Martins et al., 2012), PP-ClinTox for toxicity to human body (Gayvert et al., 2016), PP-HIV for HIV replication inhibition (hiv, ), and PP-SIDER for side effects of drugs (Kuhn et al., 2015). These involved properties are crucial especially for drug development.
Table 1 shows the statistics of SMolInstruct. It contains 3.4M samples. Each sample is a query-response pair, where the query describes a task and any task-specific information (e.g., input molecule, textual description, etc.), and the response is a sentence containing the answer to the queried task. For all the tasks, unless explicitly defined in the tasks (NC-I2F, NC-I2S, NC-S2F, and NC-S2I), we use SMILES as the representation for molecules.
4.2 SMolInstruct Construction Details
We construct the SMolInstruct dataset by following a four-step pipeline: data collection, quality control, data splitting, and instruction construction. Detailed explanations are as follows.
Data Collection. After consulting domain experts and carefully pinpointing the set of meaningful tasks (summarized in Section 4.1), we collect data for these tasks from various sources, as listed in Table 1. Specifically, for the name conversion tasks (NC-I2F, NC-I2S, NC-S2F, and NC-S2I), we leverage PubChem222https://pubchem.ncbi.nlm.nih.gov/ (Kim et al., 2019), one of the most comprehensive molecule databases. Within this database, we randomly select a large set of molecule entries, and extract their IUPAC names, SMILES representations, and molecular formulas. This obtained data is then re-organized as input-output pairs for the tasks. For molecular description-related tasks (MC and MG), we utilize a combination of ChEBI-20 (Edwards et al., 2021, 2022) and Mol-Instructions (Fang et al., 2023), as they both contain high-quality molecule-text paired data. For property prediction tasks (PP-ESOL, PP-Lipo, PP-BBBP, PP-ClinTox, PP-HIV, and PP-SIDER), we employ the well-established MoleculeNet datasets (Wu et al., 2018). We carefully select 6 datasets from MoleculeNet that represent the most essential properties for small molecule chemistry. For chemical reaction tasks (FS and RS), we collect the reaction data from USPTO-full (Lowe, 2017), which is an extensive collection encompassing over 1M reaction samples extracted from U.S. patents. All the aforementioned datasets are also widely used in previous studies (He et al., 2021; Zhong et al., 2022; Edwards et al., 2022; Irwin et al., 2022; Chen et al., 2023; Zhou et al., 2023).
Quality Control. To guarantee high quality, we apply rigorous scrutiny. The collected data contains many problematic and low-quality samples, which can be roughly categorized into the following three types, along with our curation methods: (1) Chemically invalid SMILES. Numerous SMILES strings are chemically invalid (e.g., deviating from the SMILES grammar, or violating chemical valence). To address this issue, we employ RDKit (rdk, ), a widely used toolkit for cheminformatics, to parse molecules and detect errors. (2) Wrong or inaccurate information. Based on manual check, we observed wrong and inaccurate information recorded in the data. For instance, within the USPTO-full dataset (Lowe, 2017), we identify and correct mislabeled reactants and reagents in chemical reactions by comparing their atom mappings with products. For the MC and MG tasks, we filter out those textual descriptions that lack pertinent, molecule-specific information, with a set of rules based on wording patterns, lengths and keywords. For PP-SIDER, we eliminate disorders with ambiguous names that could impede the creation of precise and comprehensible instructions. (3) Duplicated samples. They prevail in the data, and we carefully detect and remove them.
Data Splitting. Data splitting for multi-task datasets requires careful handling in order to avoid data leakage across tasks. For instance, FS and RS are reverse tasks, so data leakage occurs when the training set contains an FS sample for a certain chemical reaction and the test set has an RS sample for the same reaction. This can lead to biased evaluation. Therefore, we meticulously identify sample pairs across related tasks (FS and RS, MC and MG, and the four NC tasks) that correspond to the same molecules/reactions, and ensure that matched samples are placed together in either training or evaluation set. Moreover, some samples may share the same input but have different outputs. For instance, in the RS task, one product (the same input) may be synthesized from multiple sets of reactants (different outputs). If these samples are placed into both training and test set, it may lead to exaggerated results. Therefore we ensure that samples with identical inputs are placed together either in or outside of the test set. Additionally, to achieve fair comparisons with Mol-instructions (Fang et al., 2023), for tasks shared between the two datasets (MC, MG, FS, and RS), we ensure that their training examples are not included in our test set, allowing for a direct evaluation of their models on our test set. Following these necessary limitations, samples are randomly spitted into training/validation/test set, except for PP task samples that undergo a scaffold splitting following the canonical method (Wu et al., 2018). Table 1 shows the statistics for each split.
4.3 Merits of SMolInstruct
(1) Large-Scale. SMolInstruct consists of 3.4M distinct samples and 1.6M distinct molecules, with a diverse range of sizes, structures, and properties (see Figure 5 in Appendix C), showcasing an extensive coverage of diverse chemical knowledge.
(2) Comprehensive. SMolInstruct contains 4 types of chemical tasks (14 tasks in total), emerging as the most comprehensive instruction tuning dataset for small molecules. Notably, the tasks are meticulously selected to build a strong chemistry foundation.
(3) High-Quality. Rigorous processing steps have been implemented to exclude problematic and low-quality samples. Along with careful data splitting and canonicalization of SMILES representations SMolInstruct stands as a high-quality resource valuable for future research.
5 Experiments
5.1 Our LlaSMol Models
By fine-tuning base models on the proposed SMolInstruct dataset, we create LLMs capable of performing chemistry tasks, which we name LlaSMol (Large language models on small molecules). Specifically, we extensively consider four different LLMs as our base models, namely Galactica 6.7B (Taylor et al., 2022), Llama 2 (Touvron et al., 2023b) 7B, Code Llama (Roziere et al., 2023) 7B, and Mistral (Jiang et al., 2023) 7B, where Galactica is trained for scientific applications and has already been exposed to chemistry-related data during its pretraining, Llama 2 and Mistral are general-purpose LLMs, while Code Llama is based on Llama 2 and trained for code. We conduct instruction tuning on the proposed SMolInstruct dataset, and name the resulting models as LlaSMol, LlaSMol, LlaSMol, and LlaSMol, respectively. All the LlaSMol models are trained with LoRA (Hu et al., 2022), which is applied to all weight matrices in the self-attention and feedforward neural network (FFN) modules with lora_r and lora_alpha set to 16. The fine-tuning process utilizes the Huggingface Transformers library (Wolf et al., 2020). Training spans three epochs, employing the 8-bit AdamW optimizer, a learning rate of 1e-4, and a cosine scheduler. The input length for training is set to 512, which covers 99.7% of the samples. During inference, we adopt beam search as the generation strategy for simplicity. Details can be found in Appendix A.
5.2 Experimental Setup
(1) LLMs without fine-tuning on SMolInstruct. This type includes our four base models, namely Galactica (Taylor et al., 2022), Llama 2 (Touvron et al., 2023b), Code Llama (Roziere et al., 2023), Mistral (Jiang et al., 2023). we also benchmark against GPT-4 (Achiam et al., 2023), the current state-of-the-art (SoTA) LLM333Due to resource limitations, we evaluate GPT-4 on 500 test samples for each task, if applicable.. For Llama 2, Code Llama, and Mistral, we use 1-shot, due to their poor instruction following ability; for GPT-4, we report its results under a zero-shot setting, as GPT-4 performs best on this setting in our experiments (Appendix B). We also include Molinst (Fang et al., 2023), a Llama 2 model fine-tuned on the Mol-Instructions dataset (Fang et al., 2023), which shares the tasks of MC, MG, FS, and RS with SMolInstruct.
(2) SoTA task-specific models. To provide a comprehensive view of LlaSMol’s performance, we present results from SoTA task-specific models. For NC-I2S and NC-S2I, we compare with STOUT (Rajan et al., 2021), an encoder-decoder model trained on SMILES-IUPAC name paired data. For NC-S2F, a task achievable with a fixed algorithm, we implement a program with RDKit (rdk, ), a widely used Python toolkit for cheminformatics, and report its results. For NC-I2F where no dedicated models exist, we construct a baseline called STOUT+RDKit by aggregating STOUT for I2S conversion and RDKit for S2F conversion. For the PP tasks, our compared model is Uni-Mol (Zhou et al., 2023). It incorporates molecular 3D representations and follows a pretraining and fine-tuning paradigm. Following its original settings, we fine-tune the model on our SMolInstruct dataset with its pretrained checkpoint. In the case of MC and MG, we compare with MolT5 (Edwards et al., 2022) and directly use their released checkpoint. The reasons why we do not use our re-trained model are: (1) we were unable to reproduce results close to those reported in the paper as no original code was provided; and (2) we take great care to ensure that our test set is devoid of training examples used by MolT5, ensuring fairness in the evaluation. Lastly, regarding FS and RS, we re-train RSMILES (Zhong et al., 2022) and Molecular Transformer (Schwaller et al., 2019) for the two tasks, respectively, following their reported settings. Both of the models are transformer encoder-decoder models (Vaswani et al., 2017), specifically adapted for the FS and RS tasks.
Evaluation Metrics. We employ evaluation metrics commonly used in previous work (Schwaller et al., 2019; Zhong et al., 2022; Fang et al., 2023; Zhou et al., 2023; Chen et al., 2023), which include: (1) Exact Match (EM). This metric indicates the proportion of predicted results that exactly match the gold standards in the dataset. Each task may adopt a different definition of exact match, which can be found in Appendix A. (2) Fingerprint-based Tanimoto Similarity (FTS). This metric quantifies structural similarities between molecules by calculating the Tanimoto similarities between their Morgan fingerprints (Morgan, 1965). (3) METEOR score. For the MC task, we report the commonly-used text-based metric called METEOR score that offers a comprehensive evaluation by considering both exact matches and semantic similarity. (Lavie & Agarwal, 2007) (4) Root Mean Square Error (RMSE). Specifically used for the PP-ESOL and PP-Lipo task, this metric quantifies the accuracy of predicted values by measuring the square root of the average squared differences between predicted and actual values. (5) Accuracy (Acc). Applied in the binary classification tasks (PP-BBBP, PP-ClinTox, PP-HIV, and PP-SIDER), this metric is the ratio of correct predictions. (6) Validity (Valid). For tasks where the outputs are SMILES representations (NC-I2S, MG, FS, and RS), this metric presents the ratio of valid predictions that follow the grammar of SMILES and satisfy chemical valence rules. For all the above metrics except RMSE, the larger, the better.
For more details regarding training, inference, and evaluation setups, please refer to Appendix A.
5.3 Main Results
[b] Model NC PP I2F I2S S2F S2I ESOL Lipo BBBP Clintox HIV SIDER EM EM Valid EM EM RMSE RMSE Acc Acc Acc Acc Task-Specific, Non-LLM Based Models SoTA Existing LLMs without fine-tuning on SMolInstruct GPT-4 Galactica Lama 2 Code Llama Mistral Molinst (instruction-tuned) Our LlaSMol Series LlaSMol LlaSMol LlaSMol LlaSMol
| Model | MC | MG | FS | RS | ||||||
| METEOR | EM | FTS | Valid | EM | FTS | Valid | EM | FTS | Valid | |
| Task-Specific, Non-LLM Based Models | ||||||||||
| SoTA | ||||||||||
| Existing LLMs Without Fine-Tuning on SMolInstruct | ||||||||||
| GPT-4 | ||||||||||
| Galactica | ||||||||||
| Llama 2 | ||||||||||
| Code Llama | ||||||||||
| Mistral | ||||||||||
| Molinst (instruction-tuned) | ||||||||||
| Our LlaSMol Series | ||||||||||
| LlaSMol | ||||||||||
| LlaSMol | ||||||||||
| LlaSMol | ||||||||||
| LlaSMol | ||||||||||
Table 2 and 3 show the performance of different models on all tasks in SMolInstruct. We make the following key observations:
(1) Among all the LLMs, our LlaSMol models demonstrate the best performance on all tasks, underscoring the effectiveness of the proposed SMolInstruct dataset and the benefits of fine-tuning. When compared to our base models (Galactica, Llama 2, Code Llama, and Mistral), LlaSMol models exhibit remarkable performance enhancement after fine-tuning on SMolInstruct. This highlights SMolInstruct’s effectiveness in enhancing models’ understanding towards molecular representations and task-related knowledge, and signifies the effective learning of chemistry-related tasks by LLMs. Also, LlaSMol outperforms GPT-4 on all the tasks, despite GPT-4’s larger parameter size. Notably, LlaSMol, which has the same base model and LoRA setting as Molinst (Fang et al., 2023), surpasses it even on the shared training tasks (MC, MG, FS, and RS). It shows the benefits of our expansive and high-quality dataset compared to Mol-Instructions.
(2) Our four LlaSMol models show substantial differences in their performance, emphasizing the significant impact of base models on downstream tasks. Despite sharing identical training and inference settings, as well as comparable model sizes, LlaSMol consistently outperforms LlaSMol by a substantial margin, highlighting Mistral’s great potential on chemistry tasks. In addition, LlaSMol exhibits better performance than LlaSMol, indicating a potential synergy between programming language knowledge in Code Llama and molecular representations. For instance, understanding code grammar may have positively impacted the learning of SMILES representation, which can be regarded as a well-defined coding language for molecules. Furthermore, LlaSMol outperforms LlaSMol, and LlaSMol in most cases, suggesting the benefits of pretraining on chemistry-related documents.
(3) Our LlaSMol models exhibit comparable performance to SoTA models, showing great potential as a chemistry foundation model. Even though LlaSMol models may not outperform these SoTA models on many tasks, the observed performance gap has markedly diminished compared to previous efforts (Fang et al., 2023). Notably, with only a small proportion of parameters tuned (approximately 40M, 0.59%) in the 7B models, LlaSMol already achieves performance approaching or even surpassing the SoTA on several tasks such as NC-I2S, PP-SIDER, and MC. As further exploration in Section 5.4 will show that adding trainable parameters can lead to substantial performance boost, we anticipate that LLMs possess immense potential to surpass task-specific models through more extensive fine-tuning. Additionally, LlaSMol underscores the potential of a universal model capable of addressing multiple chemistry tasks.
For detailed experimental results on more metrics, please refer to Appendix B.
5.4 Influence of LoRA Modules and Trainable Parameters
In this section, we investigate the influence of using different LoRA modules or different sizes of trainable parameters. We take LlaSMol as the basic setting and refer to it as LlaSMol in this section for simplicity. All the compared models are listed as follows, with the trainable parameter sizes and the ratios labeled in brackets:
LlaSMol Lite (8.4M, 0.12%): LoRA is applied on q_proj and v_proj of the attention modules.
LlaSMol Attn (16.8M, 0.25%): LoRA is applied on all the attention projection matrices (including q_proj, k_proj, v_proj, o_proj).
LlaSMol FFN (23.2M, 0.34%): LoRA is applied on all the FFN projection matrices (including gate_proj, down_proj, up_proj).
LlaSMol (40.0M, 0.59%): The basic setting. LoRA is applied on all the attention and FFN projection matrices.
LlaSMol Plus (171.0M, 2.48%): LoRA is applied on all the attention and FFN projection matrices, and lm_head is set trainable.
LlaSMol Large (62.6M, 0.48%): Unlike other models that is based on Llama 2 7B, this model takes Llama 2 13B as the base model. LoRA is applied on all the attention and FFN projection matrices, same as the basic setting.
We show the model performance on the forward synthesis (FS) task in Figure 2, and the results for other tasks can be found in Appendix B. The key observations are summarized as follows: (1) Progressing from LlaSMol Lite, LlaSMol Attn, LlaSMol FFN to LlaSMol, the incorporation of more LoRA modules during training leads to a significant performance enhancement. (2) When comparing LlaSMol and LlaSMol Large, we can tell that larger base models yield superior results. Furthermore, comparing LlaSMol Large and LlaSMol Plus, despite the considerably smaller trainable parameter size, the former still outperforms the latter. This indicates that performance is not solely determined by the trainable parameter size; rather, the inherent capability of the base model also plays a crucial role. In summary, refining the selection of LoRA modules to train and employing a larger, more sophisticated base model offers significant potential for further improvements in the performance of LLMs on chemistry tasks.
6 Conclusion
To improve the performance of LLMs for chemistry, this paper introduces a large-scale, comprehensive, high-quality instruction tuning benchmark, SMolInstruct. It comprises 14 meticulously chosen tasks that are highly relevant to real-world applications and over 3M carefully curated samples that undergo thorough processing to ensure exceptional quality. Based on SMolInstruct, we developed a series of LLMs named LlaSMol on four different base models, among which we find the mistry-based LlaSMol achieves the best performance. Our comprehensive experimental results demonstrate LlaSMol’s superiority over existing LLMs including GPT-4 on all the tasks as well as substantial potential for LLMs to surpass the SoTA task-specific models. We will release our code and data to facilitate future research and development.
Impact Statements
We anticipate that our proposed SMolInstruct dataset and models will contribute to future research on LLMs for chemistry. However, we acknowledge certain limitations of this work. Firstly, despite our best efforts to ensure the high quality of the SMolInstruct dataset, we cannot guarantee the absence of incorrect or harmful information. Secondly, our primary focus is on advancing LLMs for chemistry tasks; hence we have not evaluated the models’ generalization abilities beyond those tasks or the models’ safety risks (such as generating harmful content under adversarial attacks). While recognizing the importance of such issues, we leave this topic as future work.
Appendix A Details of Experimental Setup

### v4

@@ Appendix A Preliminaries
Molecules can be represented in multiple ways, such as SMILES strings, IUPAC names, and molecular formulas. SMILES strings use a sequence of symbols to encode the 2D structures of molecules (Weininger, 1988). A molecule can have multiple SMILES strings; a canonical SMILES for the molecule is unique and deterministic. For example, the canonical SMILES representation of glucose is “C(C1C(C(C(C(O1)O)O)O)O)O”. SELFIES (Krenn et al., 2019) is an alternative representation to SMILES that also uses a sequence of symbols to denote molecular structures. Its key advantage is robustness, as every SELFIES string is guaranteed to correspond to a valid molecule. The SELFIES representation corresponding to the above SMILES representation of glucose is “[C][Branch2][Ring1][Branch1][C][C][Branch1][S][C][Branch1][N][C][Branch1][Branch2] [C][Branch1][Ring2][O][Ring1][=Branch1][O][O][O][O][O]”. Molecular formulas represent a molecule by enumerating the type and number of atoms in the molecule (Solomons et al., 2022). For example, the molecular formula for glucose is “”. IUPAC names are formal names based on natural language elements, which follow the systematic rules set by the International Union of Preferred and Applied Chemistry (IUPAC) (Favre & Powell, 2014). These names are derived from the structures and functional groups of molecules, and are intended to be human-readable. For example, the IUPAC name for glucose is “(3R,4S,5S,6R)-6-(hydroxymethyl)oxane-2,3,4,5-tetrol”.
@@
Appendix B Details of SMolInstruct
In this section, we introduce the details of our proposed dataset SMolInstruct, including statistics and construction details.
B.1 The statistics of SMolInstruct
@@
Table 5 shows the statistics of SMolInstruct. It contains 4 types of altogether 14 tasks, which are selected to be meaningful and useful. There are about 3.3M samples, and each of them is a distinct sample. In other words, there does not exist a pair of samples who share the same chemical information (i.e., the core input and output information, such as input molecules and output molecules), but with the same or different natural language templates (i.e., the task description in the query and the sentence templates in the response). When needed, one can easily create more instruction tuning samples by combining one piece of chemical information with multiple natural language templates. All in all, SMolInstruct can serve as a good benchmark for training and evaluating LLMs on various chemistry tasks.
To know more about the diversity of SMolInstruct, we conduct a statistics on the molecules. Altogether, there exist 1.6M distinct molecules, and several important statistical values are shown in Figure 2. Specifically, Bertz complexity is a topological index that measures the complexity of molecules based on the number and types of bonds and atoms. Atom count shows the number of atoms in a molecule, and it represents the size of a molecule. Molecular weight is the sum of the atomic weights of the atoms in a molecule. And ring count shows the number of rings in the molecular structures. As we can see, the values varies much, showing a extensive coverage in terms of complexity, size, and structure. Notably, when compared to Mol-Instructions (Fang et al., 2023), molecules in SMolInstruct show a higher complexity and diversity, which indicates that SMolInstruct is more comprehensive and complicated than Mol-Instructions. The scale, complexity, and diversity of SMolInstruct makes it well-suited for learning chemistry LLMs.
B.2 Details of Dataset Construction
Dataset construction involves four key steps (Section 3.2): data collection, quality control, data splitting, and instruction creation. This section provides task-specific details, omitting the common steps of canonicalizing SMILES/SELFIES and verbalizing information into query and response sentences, which have been introduced in Section 3.2.
Name Conversion (NC). The raw data for name conversion is collected from PubChem (Kim et al., 2019). Approximately 300k molecule/compound entries are randomly selected from the database, and their SMILES, IUPAC names, and molecular formulas are extracted. Entries with incomplete or missing information in these three domains are discarded. Finally, the SMILES, IUPAC names, molecular formulas are paired to create samples for the four name conversion tasks.
Property Prediction (PP). The raw data for property prediction is sourced from MoleculeNet (Wu et al., 2018). Out of its 16 core datasets555https://moleculenet.org/datasets-1, we select 6 that are only related to small molecules and are useful especially in drug discovery. Answers for regression tasks (e.g., ESOL and Lipo) are formulated as strings of numbers, and answers for binary classification tasks (e.g., BBBP and SIDER) are formulated as “Yes” or “No”.
Molecule Captioning (MC) and Molecule Generation (MG). The raw data is collected from ChEBI-20 (Edwards et al., 2021; 2022) and Mol-Instructions (Fang et al., 2023). Despite the large number of samples in Mol-Instructions, many are found to be of low quality. For example, numerous molecular descriptions end with the ambiguous phrase “with data available”, while others are overly general, making it difficult to generate a specific molecule based on the description. To ensure data quality, regular expressions and heuristic rules are employed to filter out low-quality samples.
Forward Synthesis (FS). USPTO-full (Lowe, 2017), one of the most comprehensive chemical reaction datasets, serve as the data source. The following processing steps are performed to clean the data: (1) Reactants and reagents are combined as input, and the product(s) serve as output, consistent with other datasets such as Mol-Instructions (Fang et al., 2023). (2) Duplicate chemicals in both input and output are removed to avoid redundancy. (3) If a chemical appears in both input and output, it is removed from the output to maintain data integrity. (4) Products in the output containing fewer than 5 molecules are considered non-main products and excluded. (5) If the above steps result in an empty output, the entire sample is discarded.
Retrosynthesis (RS). The data is also sourced from USPTO-full (Lowe, 2017), with the product as input and the reactants (excluding reagents) as output. During data exploration, we observe instances where reactants are mislabeled as reagents and vice versa. To address this issue, we compare the atom mapping numbers of the reactants and reagents with the products and relabel them accordingly. Subsequently, we apply the following processing steps: (1) Duplicate chemicals in both input and output are removed. (2) If a chemical appears in both input and output, it is removed from the input. (3) Products in the input containing fewer than 5 molecules are excluded. (4) In cases where multiple products exist in the input, the reaction is split into multiple samples, with each product serving as the input once. (5) If the above steps result in an empty input, the entire sample is discarded.
For all the tasks, samples containing invalid SMILES strings (i.e., those that cannot be parsed into a valid molecule with RDKit(RDKit, 2023)) are discarded, and duplicate samples are removed to avoid redundancy. Finally, since some molecules contain multiple components and they are separated by dots in SMILES, which is the same delimiter used to separate different reactants/reagents/products in FS and RS, the dots in SMILES strings for NC, PP, MC, and MG are replaced with semicolons to differentiate between these two usages.
Appendix C Comparison with Mol-Instructions
In this section, we present a comprehensive comparison between our work and Mol-Instructions (Fang et al., 2023).
We begin by comparing our dataset, SMolInstruct, with the Mol-Instructions dataset. While Mol-Instructions covers a broader scope (including molecule-oriented, protein-oriented, and biomolecular text instructions), SMolInstruct focuses exclusively on small molecules, providing a deeper and more comprehensive exploration of this domain.
If focusing on the molecule-related data, as shown in Appendix C, SMolInstruct is a larger, more comprehensive, and higher-quality dataset. It incorporates more tasks, samples, and molecular representations, and involves more careful curation. Both datasets share the tasks of MC, MG, FS, and RS. Although SMolInstruct has fewer samples for MC and MG, the included samples are of higher quality (see Section B.2). Furthermore, SMolInstruct contains substantially more samples for FS and RS, which have been carefully cleaned and processed. Additionally, SMolInstruct incorporates four NC tasks to facilitate the understanding of various molecular representations. Unlike Mol-Instructions, we do not include the reagent prediction task mainly due to the lack of sufficient high-quality data (Andronov et al., 2023) and the limited practicality of this task in real world applications.
Beyond the dataset, our work makes contributions to the exploration of chemistry LLMs. While Fang et al. (2023) primarily focus on the dataset itself and provide a preliminary exploration of the models, we conduct comprehensive experiments to investigate the abilities of LLMs in the chemistry domain. Our experiments in Section 4 demonstrates that our LlaSMol models achieves superior performance compared to the LLMs trained on Mol-Instructions and the strongest LLMs such as GPT-4 and Claude 3 Opus, greatly diminishing the gap between LLMs and SoTA task-specific models. Moreover, we provides valuable insights about multi-task training, LoRA (Hu et al., 2022) settings, and other aspects that could be helpful for future research in this field.
| Mol-Instructions | SMolInstruct (ours) | ||
| Tasks | name conversion | [X] No this task. | 1.2M samples |
| property prediction | 362.1k samples on HOMO/LUMO energy. | 78.3k samples on 6 useful properties. | |
| molecule captioning | 298.3k samples. | 60.3k samples. | |
| molecule generation | 298.3k samples. | 60.3k samples. | |
| forward synthesis | 125.4k samples. | 977.9k samples. | |
| retrosynthesis | 129.7k samples. | 948.0k samples. | |
| reagent prediction | 125.4k samples. | [X] Not included due to its insufficient data and limited practicality. | |
[…]
| #Samples | 1.3M | 3.3M | |
[…]
Appendix D Details of Experimental Setup